# BoostTrack++: using tracklet information to detect more objects in multiple object tracking

### v1

@@ 1 Introduction
Recently, BoostTrack  [14] used a one-stage association combined with strategies to increase (boost) the detection confidence of some detected bounding boxes before discarding detections with low confidence scores. In confidence boost based on detection of likely objects (DLO), it used intersection over union (IoU) between detected bounding boxes and the tracklets to discover the low-confidence detections with high overlap with existing tracklet. BoostTrack boosted the confidence score of these detections assuming they correspond to currently tracked objets, which improved the tracking performance in crowded scenes with frequent occlusions (namely, on the MOT20 dataset). However, the method resulted in an increased number of new IDs, and identity switches, which indicates a more sophisticated method is required. In this paper, we extend the idea of DLO confidence boost.
@@
Another weakness of the BoostTrack DLO confidence boost technique is that it treats all tracklets equally when deciding if IoU between a tracklet and a detection is high enough. However, the tracklets that are not recently updated usually have relatively low IoU with the corresponding detections once they are matched. Setting a threshold for "high enough" similarity should be tracklet specific and depend on the number of frames since the last time the tracklet was updated. For example, if a tracklet was not matched for 30 frames, IoU of 0.8 can be considered very high. We attempt to solve this problem by introducing varying threshold based on the number of frames since the last update of a given tracklet.
@@ 2.1 Tracking by detection
[…]
@@
where $u$, $v$, $h$ and $r$ represent the coordinates of the object center, the height of the object and the ratio of the object’s width and height, respectively. By $\dot{u},\dot{v},\dot{h},\dot{r}$ we denote their corresponding velocities. When a tracklet is associated with detected bounding box $D_{i}$, that bounding box is used is used as the observation in Kalman update step.
@@ 2.3 BoostTrack
Adding appearance similarity for the association to BoostTrack results in the BoostTrack+ method.
@@ 2.4 Buffered IoU
To account for irregular motions, Buffered IoU (BIoU) is introduced in  [15]. If the predicted state is inaccurate (due to irregular motion), the IoU between the predicted bounding box and the detected bounding box will be low (possibly 0) which makes association difficult or even impossible. The authors proposed to scale (add "buffers" to) the detected and the predicted bounding boxes. Let $o=(x,y,w,h)$ be the original detection (or tracklet), where $(x,y)$ represents the top-left coordinate of the bounding box, and $w$ and $h$ its width and height, respectively. Authors propose to use scaled detection $o_{b}=(x-bw,y-bh,w+2bw,w+2bh)$, where $b\geq 0$ is the scale parameter. More specifically, they performed two-stage association: in the first stage they used small scale parameter $b_{1}$, and in the second stage they associated remaining tracklets and detections using larger scale parameter $b_{2}$. The logic behind two-stage association is that the unmatched tracklets are more difficult to match and require larger bounding boxes for successful matching.
@@ 3 Limitations of confidence boost based on detection of likely objects
One of the contributions of BoostTrack  [14] is the DLO confidence boost technique. The effectiveness of this technique is most notable in the case of the MOT20 dataset, where the DLO boost increased MOTA [^2] We discuss the metrics used in subsection  5.1. score by 4.8% (see table  1).
@@
The most important issue is the introduction of new IDs, which should not happen if the detections with boosted confidence truly correspond to existing tracklets. Ideally, a DLO confidence boost should produce no new IDs or result in identity switches (IDSWs). This issue is best illustrated by results on the MOT20 validation set which we present in table  1. On the MOT17 validation set (which has only 339 ground-truth IDs, as opposed to the MOT20 validation set which contains 1418 IDs) the DLO boost is less significant.
@@
If we take SORT  [31] as the most basic and representative baseline, we notice that the increase in main MOT metrics (the left side of table  1), comes at the price of increased IDs (+8%) and IDSWs (+12%).
@@
Several papers have discussed the limitations of IoU used as a similarity measure for association and used various IoU modifications or other motion cues to construct a richer similarity measure (e.g.  [25, 32, 33, 30, 14, 34]). If using IoU alone can cause ambiguities and IDSWs in association, it is also unreliable to base DLO confidence boost solely on IoU. If a richer similarity measure improves association, it should also improve the DLO confidence boost.
@@
First, it does not take into account the original confidence score. A detection with a higher confidence score is more likely to be a true positive. If a detection has the confidence score slightly below the threshold $\tau$, it should require only a mild conformation in similarity with existing tracklet to boost its confidence above the $\tau$.
@@
As in BoostTrack  [14], we use Kalman filter as the tracking module. The quality of Kalman filter predictions reduces when the tracklet does not get matched with a detection, i.e. when we miss the observation and do not execute the Kalman update step. Error covariance prediction in step $t$, $\hat{P}_{t}$, is calculated as $\hat{P}_{t}=F\cdot P_{t-1}F^{\mathsf{T}}+Q$ (we provide the definition of the state $x$ and state transition matrix $F$ in equation  2).
The estimate of variance of $x\in\{u,v,h,r\}$, after not executing Kalman update for $n$ steps (frames), if we omit $Q$ for simplicity, becomes:
@@
The variance of the predictions increases quadratically with the increase of $n$ [^3] Equality  5 follows directly from substituting $F$ and $x$ from equation  2 and applying prediction for $n$ steps..
Figure 2: IoU statistics calculated on MOT17 and MOT20 validation sets for tracklets with different nummber of steps since last update. a) Average IoU b) 0.9 quantile of IoU values.
We empirically verified that the IoU value between a detection and a tracklet decreases as the number of steps since the last successful match (last position update) of the given tracklet increases. The left side of the figure  2 shows the average IoU value (calculated on MOT17  [16] and  [17] validation sets) depending on the number of frames since the last update, while the right side of the same figure shows values of $0.9$ quantiles. The $0.9$ quantile gives us information about what can be considered "high" IoU depending on the number of steps since the last update of the corresponding tracklet. As the number of steps since the last update increases, we can observe a decrease in average IoU, and, more importantly, decrease in $0.9$ quantile.
@@
The data from table  1 and the discussion presented indicate that a more sophisticated DLO confidence boost is needed, which we attempt to offer in the following section.
@@ 4.1 Soft BIoU
Every predicted bounding box should be enlarged to the extent which is proportional to the uncertainty in the quality of the prediction. In  [14], tracklet confidence is defined and used as a measure of reliability of the predicted tracklet state. We propose to use tracklet confidence to calculate tracklet specific scale without the need for a two-stage association. However, if we use different scales for different tracklets and comparing with detections in one stage, we cannot simply scale the traclets and detections by the same parameter $b$ and calculate IoU between all pairs, as done in  [15]. Every detection-tracklet pair $(D_{i},T_{j})$ should use a specific scale based on tracklet confidence, $c_{t_{j}}$, value.
@@
where $(x,y),w$, and $h$ represent top-left corner of the bouding box, its width and height, respectively. We define soft BioU (SBioU) between $D_{i}$ and $T_{j}$ as
@@
for $i\in\{1,2,\ldots,n\}$, $j\in\{1,2,\ldots,m\}$. As the tracklet confidence decreases, the scale increases. To calculate the scaling parameter, we divide $1-c_{t_{j}}$ by $2$ to match the range of scale values from  [15] (for $c_{t_{j}}=0.0$ it increases scale to $b_{2}=0.5$). When scaling the detection box, we use smaller scale since we are more uncertain about the tracklet state compared to the detected bounding box position. Note that SBIoU reduces to IoU when tracklet confidence is equal to 1.
@@ 4.2 Using improved similarity measure to find likely objects
DLO confidence boost relies on IoU similarity. We propose to replace IoU in equation  4 with more sophisticated similarity measure $S$, which gives:
@@
Trivially, we could set $S=\text{SBIoU}$. However, we propose to use a richer similarity measure, of which soft BIoU is only a part of.
In  [14], a combination of IoU, shape and Mahalanobis distance similarity was used to construct a "boosted" similarity measure which improved association performance. If such similarity measure helps to better distinguish objects, it should also improve performance of detecting the likely objects.
@@
Intuitively, using average of multiple similarity measures for the DLO confidence boost means that all similarity measures need to "agree" to increase the confidence score of a given detection. In our implementation, for the DLO confidence boost we use:
@@
On the left side of the figure we show false positive detection corresponding to a ghost track. The IoU between the detection and the tracklet is high enough (IoU=0.82). However, the ghost tracks are static and the corresponding estimated covariance values are low which leads to low Mahalanobis distance similarity values even for a small mismatch ($S^{MhD}=0.03$ in the example). This results in low average similarity and discarding the false positive detection.
On the right side of the figure we show a true positive detection $D_{i}$ ($c_{d_{i}}=0.16$) with relatively low IoU between the corresponding tracklet $T_{j}$ ($\text{IoU}(D_{i},T_{j})=0.74$). However, $\text{SBIoU}(D_{i},T_{j})=0.77,S^{MhD}(D_{i},T_{j})=0.96$, $S^{shape}(D_{i},T_{j})=0.82$, and the average similarity (from equation  9) is high enough to keep $D_{i}$ for the association step.
@@ 4.3 Soft detection confidence boost
There is a positive correlation between the IoU (calculated between the ground-truth and detected bounding boxes) and the detection confidence score  [35]. If the tracking module gives accurate predictions, the IoU between a tracklet and the ground-truth should be high and the same correlation applies between the tracklet-detection IoU and the detection confidence. The IoU between the tracklet and the detection can thus be a measure or an indicator of detection confidence. Following the described logic, for every detected bounding box $D_{i}$, $D_{i}\in\{D_{1},D_{2},\ldots,D_{n}\}$, we perform soft detection confidence boost to obtain new detection confidence score $\hat{c}_{d_{i}}$ as:
@@
where $\alpha\in[0,1]$ and $q\geq 1$ are hyperparameters and $S$ is any similarity measure. We raise $S$ to the power $q$ to have a better control over the entire process (all hyperparameter values are discussed in subsection  5.2). Equation  11 combines original confidence scores with the similarity and solves the problem of equal treatment of low and relatively high confidence score detections.
@@
Figure  4 shows a frame from MOT17 (sequence 11) and a detected bounding box (in red) with confidence a score 0.56 which is slightly below the standard threshold 0.6 used for MOT17. The IoU between the detection and the corresponding tracklet is relatively low (0.82). However, since the original confidence score is only slightly below the threshold, the imperfect IoU of 0.82 is enough to boost the detection confidence above the threshold.
@@ 4.4 Varying similarity threshold
All the similarity measures discussed previously depend upon tracking performance, i.e. the quality of predictions. As noted in section  3, the quality of predictions reduces with the increase of a number of steps since the last tracklet update. Reduced quality of predictions leads to lower IoU value when objects get matched (see figure  2) and indicates that single threshold value for DLO confidence boost cannot provide the best results.
To address the described issue, we propose to use different threshold values when searching for the likely objects. The threshold should be higher for the IoU (or any other similarity measure) between a detection and recently updated tracklet.
[…]
@@ 4.5 Combining the proposed methods
All proposed additions are mutually independent and can be used separately or jointly. Algorithm  1 shows the combined usage of all proposed elements and can be used as a replacement for the DLO confidence boost used in BoostTrack  [14]. Parameters $useS$, $useSB$ and $useVT$ control whether we should use improved similarity measure, soft detection confidence boost and varying threshold, respectively.
@@
14:                  $c_{d_{i}}:=\text{max}\Big{(}c_{d_{i}},\alpha\cdot c_{d_{i}}+(1-\alpha)\cdot\big{(}\max_{j}(S(D_{i},T_{j})\big{)}^{q}\Big{)}$triangleright Applying equation  11
@@
27:     return $c_{d_{1}},c_{d_{2}},\ldots,c_{d_{n}}$triangleright Outputs boosted confidence scores $\hat{c}_{d_{1}},\hat{c}_{d_{2}},\ldots,\hat{c}_{d_{n}}.$

### v2

@@ 1 Introduction
Recently, BoostTrack  [14] used a one-stage association combined with strategies to increase (boost) the detection confidence of some detected bounding boxes before discarding detections with low confidence scores. In confidence boost based on detection of likely objects (DLO), it used intersection over union (IoU) between all detected bounding boxes and all tracklets to discover the low-confidence detections with high overlap with existing tracklet. BoostTrack boosted the confidence score of these detections assuming they correspond to currently tracked objets, which improved the tracking performance in crowded scenes with frequent occlusions (namely, on the MOT20 dataset). However, the method resulted in an increased number of new IDs, and identity switches, which indicates a more sophisticated method is required. In this paper, we extend the idea of DLO confidence boost.
@@
Another weakness of the BoostTrack DLO confidence boost technique is that it treats all tracklets equally when deciding if IoU between a tracklet and a detection is high enough. However, the tracklets that are not recently updated (i.e. not matched with a detected bounding box for multiple subsequent frames due to occlusion, detector or association failure) usually have relatively low IoU with the corresponding detections once they are matched. Setting a threshold for "high enough" similarity should be tracklet specific and depend on the number of frames since the last time the tracklet was updated. For example, if a tracklet was not matched for 30 frames, IoU of 0.8 can be considered very high. We attempt to solve this problem by introducing varying threshold based on the number of frames since the last update of a given tracklet.
@@ 2.1 Tracking by detection
$x_{k}$ $=F\cdot x_{k-1}+q_{k}\quad\text{(dynamic equation)},$ (1)
$y_{k}$ $=H\cdot x_{k}+r_{k}\quad\text{(observation equation)},$
@@
where $u$, $v$, $h$ and $r$ represent the coordinates of the object center, the height of the object and the ratio of the object’s width and height, respectively. By $\dot{u},\dot{v},\dot{h},\dot{r}$ we denote their corresponding velocities. When a tracklet is associated with detected bounding box $D_{i}$, that bounding box is used as the observation in Kalman update step.
@@ 2.3 BoostTrack
The authors also proposed BoostTrack+ method which uses appearance similarity (between embedding vectors) in addition to similarity measures used in  3. BoostTrack+ outperforms BoostTrack at the expanse of longer computation time.
In table  1, we summarize the methods used as baselines in this paper.
Table 1: Summary of various baseline methods.
Method Description Pros Cons Improvements over prev. method
SORT Tracker that uses IoU One-stage association, Poor selection of /
as similarity measure real-time performance true-positive detections,
poor tracking performance
BoostTrack Tracker focused on One-stage association, Lacks strong cues Improved overall performance
handling unreliable detections uses all detections, such as visual
and improving association real-time performance features embedding,
introduces IDs and IDSWs
BoostTrack+ Tracker focused on One-stage association, Cannot operate in Improved tracking performance
handling unreliable detections uses all detections, real time in crowded scenes,
and improving association introduces IDs and IDSWs
@@ 2.4 Buffered IoU
To account for irregular motions, Buffered IoU (BIoU) is introduced in  [15]. If the predicted state is inaccurate (due to irregular motion), the IoU between the predicted bounding box and the detected bounding box will be low (possibly 0) which makes association difficult or even impossible. The authors proposed to scale (add "buffers" to) the detected and the predicted bounding boxes, i.e. tracklets. Let $o=(x,y,w,h)$ be the original detection (or tracklet), where $(x,y)$ represents the top-left coordinate of the bounding box, and $w$ and $h$ its width and height, respectively. The authors propose to use scaled detection $o_{b}=(x-bw,y-bh,w+2bw,h+2bh)$, where $b\geq 0$ is the scale parameter. More specifically, they performed two-stage association: in the first stage they used small scale parameter $b_{1}$, and in the second stage they associated remaining tracklets and detections using larger scale parameter $b_{2}$. The logic behind two-stage association is that the unmatched tracklets are more difficult to match and require larger bounding boxes for successful matching.
@@ 3 Limitations of confidence boost based on detection of likely objects
One of the contributions of BoostTrack  [14] is the DLO confidence boost technique. The effectiveness of this technique is most notable in the case of the MOT20 dataset, where the DLO boost increased MOTA [^2] We discuss the metrics used in subsection  5.1. score by 4.8% (see table  2).
@@
The most important issue is the introduction of new IDs, which should not happen if the detections with boosted confidence truly correspond to existing tracklets. Ideally, a DLO confidence boost should produce no new IDs or result in identity switches (IDSWs). This issue is best illustrated by results on the MOT20 validation set which we present in table  2. On the MOT17 validation set (which has only 339 ground-truth IDs, as opposed to the MOT20 validation set which contains 1418 IDs) the DLO boost is less significant.
@@
If we take SORT  [31] as the most basic and representative baseline, we notice that the increase in main MOT metrics (the left side of table  2), comes at the price of increased IDs (+8%) and IDSWs (+12%).
@@
Several papers have discussed the limitations of IoU used as a similarity measure for association and used various IoU modifications or other motion cues to construct a richer similarity measure (e.g.  [25, 32, 33, 30, 14, 34]). If using IoU alone can cause ambiguities and IDSWs in association, it is also unreliable to base the DLO confidence boost solely on IoU. Using a richer similarity measure improves association and enables the discard of falsely associated detection-tracklet pairs. Since DLO confidence boost relies on calculating similarity, a richer similarity measure should also improve the selection of true positive detections.
@@
First, it does not take into account the original confidence score. A detection with a higher confidence score is more likely to be a true positive. If a detection has a confidence score slightly below the threshold $\tau$, it should require only a mild conformation in similarity with existing the tracklet to boost its confidence above the $\tau$.
@@
As in BoostTrack  [14], we use the Kalman filter as the tracking module. The quality of Kalman filter predictions reduces when the tracklet does not get matched with a detection, i.e. when we miss the observation and do not execute the Kalman update step. Error covariance prediction in step $t$, $\hat{P}_{t}$, is calculated as $\hat{P}_{t}=F\cdot P_{t-1}F^{\mathsf{T}}+Q$ (we provide the definition of the state $x$ and state transition matrix $F$ in equation  2).
The estimate of the variance of $x\in\{u,v,h,r\}$, after not executing Kalman update for $n$ steps (frames), if we omit $Q$ for simplicity, becomes:
@@
The variance of the predictions increases quadratically with the increase of nn [^3] Equality  5 follows directly from substituting $F$ and $x$ from equation  2 and applying prediction for $n$ steps..
Figure 2: IoU statistics calculated on MOT17 and MOT20 validation sets for tracklets with different numbers of steps since the last update. a) Average IoU b) 0.9 quantile of IoU values.
We empirically verified that the IoU value between a detection and a tracklet decreases as the number of steps since the last successful match (last position update) of the given tracklet increases. The left side of the figure  2 shows the average IoU value (calculated on MOT17  [16] and MOT20  [17] validation sets [^4] For more information on datasets and implementation details, see subsections 5.1 and 5.2.) depending on the number of frames since the last update, while the right side of the same figure shows values of $0.9$ quantiles. The $0.9$ quantile gives us information about what can be considered "high" IoU depending on the number of steps since the last update of the corresponding tracklet. As the number of steps since the last update increases, we can observe a decrease in average IoU, and, more importantly, a decrease in $0.9$ quantile.
@@
The data from table  2 and the discussion presented indicate that a more sophisticated DLO confidence boost is needed, which we attempt to offer in the following section.
@@ 4.1 Soft BIoU
Every predicted bounding box should be enlarged to the extent which is proportional to the uncertainty in the quality of the prediction. In  [14], tracklet confidence is defined and used as a measure of reliability of the predicted tracklet state. We propose to use tracklet confidence to calculate tracklet specific scale without the need for a two-stage association. However, if we use different scales for different tracklets and comparing with detections in one stage, we cannot simply scale the tracklets and detections by the same parameter $b$ and calculate IoU between all pairs, as done in  [15]. Every detection-tracklet pair $(D_{i},T_{j})$ should use a specific scale based on tracklet confidence, $c_{t_{j}}$, value.
@@
where $(x,y),w$, and $h$ represent top-left corner of the bounding box, its width and height, respectively. We define soft BioU (SBioU) between $D_{i}$ and $T_{j}$ as
@@
for $i\in\{1,2,\ldots,n\}$, $j\in\{1,2,\ldots,m\}$. As the tracklet confidence decreases, the scale increases. To calculate the scaling parameter, we divide $1-c_{t_{j}}$ by $2$ to match the range of scale values from  [15] (for $c_{t_{j}}=0.0$ it increases scale to $b_{2}=0.5$). When scaling the detection box, we use a smaller scale since we are more uncertain about the tracklet state compared to the detected bounding box position. Note that SBIoU reduces to IoU when tracklet confidence is equal to 1.
@@ 4.2 Using improved similarity measure to find likely objects
DLO confidence boost relies on IoU similarity. We propose to replace IoU in equation  4 with a more sophisticated similarity measure $S$, which gives:
@@
Trivially, we could set $S=\text{SBIoU}$. However, we propose to use a richer similarity measure, of which soft BIoU is only a part.
In  [14], a combination of IoU, shape and Mahalanobis distance similarity was used to construct a "boosted" similarity measure which improved association performance. If such a similarity measure helps to better distinguish objects, it should also improve the performance of detecting the likely objects.
@@
Intuitively, using an average of multiple similarity measures for the DLO confidence boost means that all similarity measures need to "agree" to increase the confidence score of a given detection. We use average for simplicity and to give all summands equal weight (note that in equation  3 weights, i.e. lambdas, do not have to be equal). In our implementation, for the DLO confidence boost we use:
@@
On the left side of the figure, we show false positive detection corresponding to a ghost track. The IoU between the detection and the tracklet is high enough (IoU=0.82). However, the ghost tracks are static and the corresponding estimated covariance values are low which leads to low Mahalanobis distance similarity values even for a small mismatch ($S^{MhD}=0.03$ in the example). This results in low average similarity and discarding the false positive detection.
On the right side of the figure, we show a true positive detection $D_{i}$ ($c_{d_{i}}=0.16$) with relatively low IoU between the corresponding tracklet $T_{j}$ ($\text{IoU}(D_{i},T_{j})=0.74$). However, $\text{SBIoU}(D_{i},T_{j})=0.77,S^{MhD}(D_{i},T_{j})=0.96$, $S^{shape}(D_{i},T_{j})=0.82$, and the average similarity (from equation  9) is high enough to keep $D_{i}$ for the association step.
@@ 4.3 Soft detection confidence boost
The greater the detection confidence score, the more accurate the corresponding detected bounding box, i.e. the greater IoU between the ground-truth and detected bounding box  [35]. Provided that the tracking module gives accurate predictions, there is also a positive correlation between the detection confidence score and the IoU between the tracklet (i.e. predicted bounding box) and the corresponding detected bounding box. IoU between the tracklet and the detection can thus be a measure or an indicator of detection confidence. If the IoU between the detected bounding box $D_{i}$, $D_{i}\in\{D_{1},D_{2},\ldots,D_{n}\}$, and some tracklet is high, the detection confidence score $c_{d_{i}}$ should also be high. Following the described logic, for every detected bounding box $D_{i}$, we perform soft detection confidence boost to obtain a new detection confidence score $\hat{c}_{d_{i}}$ as:
@@
where $\alpha\in[0,1]$ and $q\geq 1$ are hyperparameters and $S$ is any similarity measure. We raise $S$ to the power $q$ to have better control over the entire process (all hyperparameter values are discussed in subsection  5.2). Equation  11 combines original confidence scores with the similarity and solves the problem of equal treatment of low and relatively high confidence score detections.
@@
Figure  4 shows a frame from MOT17 (sequence 11) and a detected bounding box (in red) with confidence score 0.56 which is slightly below the standard threshold 0.6 used for MOT17. The IoU between the detection and the corresponding tracklet is relatively low (0.82, which is low compared to the threshold value 0.923 used for MOT17 in  [14]). However, since the original confidence score is only slightly below the threshold, the imperfect IoU of 0.82 is enough to boost the detection confidence above the threshold.
@@ 4.4 Varying similarity threshold
All the similarity measures discussed previously depend upon tracking performance, i.e. the quality of predictions. As noted in section  3, the quality of predictions reduces with the increase of the number of steps since the last tracklet update. Reduced quality of predictions leads to lower IoU value when objects get matched (see figure  2) and indicates that a single threshold value for DLO confidence boost cannot provide the best results.
To address the described issue, we propose to use different threshold values when searching for the likely objects. The threshold should be higher for the IoU (or any other similarity measure) between a detection and a recently updated tracklet.
Let $\text{last\_update}(T_{j})$ be the number of steps since the last update of tracklet $T_{j},j\in\{1,2,\ldots,m\}.$ We decrease the threshold $\beta_{j}$ corresponding to a tracklet $T_{j}$ linearly from the starting value $\beta_{high}$ to the final value $\beta_{low}$, and define our varying threshold $\beta_{j}$ as:
@@ 4.5 Combining the proposed methods
All proposed additions are mutually independent and can be used separately or jointly. Algorithm  1 shows the combined usage of all proposed elements and can be used as a replacement for the DLO confidence boost used in BoostTrack  [14]. Parameters $useS$, useSBuseSB and $useVT$ control whether we should use improved similarity measure, soft detection confidence boost and varying threshold, respectively.
@@
14:      $c_{d_{i}}:=\text{max}\Big{(}c_{d_{i}},\alpha\cdot c_{d_{i}}+(1-\alpha)\cdot\big{(}\max_{j}(S(D_{i},T_{j})\big{)}^{q}\Big{)}$ $\triangleright$ Applying equation  11
@@
27:  return $c_{d_{1}},c_{d_{2}},\ldots,c_{d_{n}}$ $\triangleright$ Outputs boosted confidence scores $\hat{c}_{d_{1}},\hat{c}_{d_{2}},\ldots,\hat{c}_{d_{n}}.$